OpenStudy (anonymous):

how do i simplify lnx^3?

OpenStudy (anonymous):

There is basics rule of logarithm: \[\ln(x)^n = n \cdot \ln(x)\]

OpenStudy (anonymous):

Are you given the value of x??

OpenStudy (anonymous):

I am expecting a reply from your side..

OpenStudy (anonymous):

im sorry. i got x^3 as my answer

OpenStudy (anonymous):

idk if its right though